# A Transformer approach for Electricity Price Forecasting

### v2

@@ Abstract
Index Terms:
Electricity Price Forecasting, Transformers, Attention, , paper, template.
I Introduction
@@
Forecasting electricity prices plays a pivotal role in the modern power systems landscape. Accurate and efficient price forecasting methods have become increasingly crucial as the electricity market continues to evolve towards a more competitive and deregulated structure  [1, 2] . These forecasts serve as a fundamental decision-making tool for various market participants, including power producers, consumers, traders, and grid operators.
@@
Furthermore, with the growing integration of renewable energy sources, which are inherently intermittent and unpredictable, the volatility and complexity of electricity prices have increased. This motivates the need for more sophisticated and robust forecasting methods to capture these complex dynamics and provide accurate price forecasts  [3]. Thus, developing effective electricity price forecasting methods is not only an economic imperative but also a key enabler for the sustainable and reliable operation of power systems.
@@
The techniques for electricity price forecasting (EPF) have evolved significantly over the years, with recent advancements in machine learning and artificial intelligence leading the way  [4]. While effective in certain scenarios, traditional statistical methods often struggle to capture the complex, non-linear dynamics of electricity prices. This has led to the exploration of more sophisticated models, particularly neural networks and deep learning models applied to time series forecasting. For example, Recurrent Neural Networks (RNNs) and Long Short Term Memory (LSTM) networks have been successfully employed for EPF  [5]. These models are capable of learning complex patterns and long-term dependencies, making them well-suited for forecasting tasks.
@@
Among the deep learning models, the application of Transformer-based architecture is an emerging technique. The design of attention mechanisms for approaching sequence-to-sequence problems started in the NLP field, where the output is not only a label, but a multiple sequence (i.e. a complete sentence). Traditionally an encoder-decoder structure was used until the Attention layer was presented [6]. The Attention layer allows the decoder to focus its attention on a specific word or group of words from the input of the encoder (the original sentence in the case of translation). This helped improve the State of the Art of many NLP problems, becoming the basis of the Transformer model. One of the advantages of the Transformer architecture is that it enables the generation of much bigger Deep Learning models (such as the BERT model [7]).
@@
The attention mechanisms that allow focus on the most relevant parts of the input make them powerful tools for prediction tasks and have been successfully used in several fields such as Autonomous Vehicles [8] or Image Classification [9].
@@
Attention mechanisms have also been applied to the power sector. For example,  [10] use attention mechanisms combined with LSTM models for short-term wind power forecasting. Other examples include [11] and [12] which propose the use of attention and transformers for electricity load forecasting.
@@
Regarding electricity price forecasting, additional challenges are presented as input explanatory variables have to be included as relevant predictors in the model, hence, [13] uses an attention layer combined with BLSTM layers for electricity price prediction to achieve promising results. Additionally, [14] propose a Transformer-based model combined with GRU layers for probabilistic electricity price forecasting.
@@
Therefore, the application of Transformer models in forecasting is still in its early stages, and further research is needed to understand their potential and limitations with respect to other forecasting methods. In fact, most of the mentioned models were not Transformers, but a recurrent model combined with attention layers. Authors such as [15], still question the effectiveness of Transformer capabilities, hence, making this a relevant issue in the research community.
@@
Additionally, in the EPF field, the issue of reproducibility has been a significant challenge. Many studies have utilized unique, non-public datasets and have tested their methods over too short and limited market samples, making it difficult to evaluate the effectiveness of new predictive algorithms. In response to these issues,  [2] proposed the EPF toolbox, an open-source toolbox for electricity price forecasting that enhances reproducibility, promotes transparency and drives innovation.
@@
II Proposed Transformer model for Electricity Price Forecasting
This section explains the proposed model for EPF. The objective is the same as stated in [2]: to predict the electricity prices of the 24h of the day $D+1$ based on the exogenous variables of the 24h of the day $D+1$ as well as past prices and exogenous variables.
The model proposed in this paper is a pure Transformer, specifically, a Transformer Encoder as in the case of BERT [16] or ViT [9]. With this structure, the Transformer can capture temporal patterns and make a prediction focusing more attention on longer trends than other temporal models (e.g. LSTMs). This is evidenced in the case of Natural Language Processing  [17], where the Transformer does not forget the previous information in longer sequences.
@@ II Proposed Transformer model for Electricity Price Forecasting
We ought to establish a direct relationship between exogenous variables and prices and keep the Transformer to model the price dynamics. Therefore, it is decided to include only past prices as historical information. Therefore, this model will have two inputs, the exogenous values for the 24h of the day $D+1$ and the prices of the 24h for the past $m$ days. The general structure can be visualized in Figure 1. Then, the interval used for the past prices will be [$D-m$, $D$], including both days. These $m$ days will be a hyperparameter of the model called sequence length that should be optimized.
@@
III Case Study
As pointed out in [2], in the literature the comparisons between different types of models have been unfair most times. They detected the following problems in the published literature on EPF models:
@@ III Case Study
Hence they proposed a new framework for testing the different models: epftoolbox [18]. This python library allows to use the models they have developed and also test the different datasets to build a fair comparison, solving all the issues that have been mentioned previously. Therefore, in this paper, this framework is used to compare and test the model presented.
III-A Open Access Datasets
To find the best algorithm for electricity price forecasting, it is important to consider multiple datasets, rather than relying on a single one. In [2], five different datasets for different countries are available so methods can be compared fairly:
@@ III-A Open Access Datasets
III-B Train-Validation-Test split
The testing period is a source of discussion. Most articles try to evaluate algorithms in short periods, e.g. a week or a month. However, the dynamics in a market can change throughout the year, especially in the electricity domain, which is highly affected by the time of the year due to issues such as the available amount of renewable energies or the time consumers are spending at home (great difference between summer and winter). Following [2], a two-year testing period is used to account the different dynamics that can appear.
@@ III-B Train-Validation-Test split
TABLE I: Train-Test split
III-C Retraining and Renormalization
In [2], the authors state that a model should be retrained instead of being directly evaluated in the whole dataset, or otherwise it is not tested in realistic conditions, since in real-life forecasting models are often retrained to account for the last market information. Even though this idea is logical, it generates two main problems for testing more advanced models such as the one proposed in this paper
On the one hand, although retraining each day simulate well real life behaviour, there are some difficulties to automatize that retraining for more complicated models. While retraining a Deep Neural Network composed of linear layers and activation functions seems to follow a straightforward process explained in [2], where early-stopping gives the optimal training point for the selected hyperparameters, more complicated models as Transformers can be more difficult to train due to issues like vanishing or exploding gradients or the amplification effect pointed put in [19]. Therefore for these models, it is more likely that a Machine Learning Engineer monitors the retraining phase and execute it if necessary. It could be the case that for some time the parameters for the retraining are the same, and only when the data changes are significant enough it is necessary to look again for optimal values.
On the other hand, the computational burden is expensive for advanced models. While the best model in [2] takes around five minutes to train, the Transformer model presented in this paper takes around twenty minutes (measured in a computer with a GPU RTX 3060).
@@ III-C Retraining and Renormalization
Nevertheless, as suggested in [2], a renormalization step was applied consisting on subtracting the mean and standard deviation from the data previous to the next day to predict. This helps the model to adapt if the conditions vary greatly during the test period.
III-D Models to Compare
@@
  • DNN: Regarding neural networks in  [2], a Deep Neural Network with two layers was selected as the best Machine Learning Model. In fact, the model that proved to be more accurate in their study was the Ensemble of Deep Neural Networks. Therefore, this model ensemble is selected as a benchmark. As it was explained the ensemble will be trained on the training period and evaluated during the test period using a renormalization step but not retraining. We compute two approaches, a renormalization with all the past data and a renormalization with only the last year values (as in [2]).
@@
  • Transformer model: This is the transformer model proposed. As mentioned in Section II, some hyperparameters have to be optimized using a validation set. Following what was proposed in [2], validation is constructed with the last 42 weeks of the training set. Some values were proposed for the different hyperparameters and datasets and the configuration with the best performance is selected. In Tables VI, VII, VIII, IX and X the results of some of the configurations in the different datasets are presented along with the best ones highlighted in gray. The best model is evaluated then on the test set, without any retraining but with data renormalization with all past values.
Following what was exposed in [2], to measure the results offered by the model presented in this paper and compare it with the benchmarks explained, the following metrics will be used: MAE, RMSE and sMAPE.
III-E Statistical Testing
@@
In particular, the one-sided DM test is implemented in [2] and used in the case study. The null hypothesis of the test is that the predictions of model A are more accurate than the ones of a model B. The lower the p-value obtained, the more observed data inconsistent with the null hypothesis. Taking into account the convention to reject the null hypothesis at p-values lower than 5%, if a lower p-value is obtained, it can be declared that the difference between predictions of model B and predictions of model A are statistically significant, being B the model with better performance.
III-F Software
The aim of this paper is to compare the Transformer model presented with the ones developed in the  [2]. Hence, every implementation from the mentioned library will be used in order to follow a rigorous approach to compare the different methods. As the aim of this paper is to openly contribute to EPF research, all the code and the best models can be found in https://github.com/osllogon/epf_transformers.
IV Results
The results of the selected models tested on the different datasets are presented in this section. For the comparison, the different metrics explained in the Section LABEL:sec:_case_study,_definition,_metrics are calculated for the test set predictions.
Tables II, III and IV show the benchmark Naïve, DNN with all past renormalization and DNN with last-year renormalization model results, respectively. As can be observed, in some cases the results of the DNN model are better and, in other cases, the naive model outperforms the DNN Ensembles. This could be due to the fact that the DNN model has to make predictions for a long test period without retraining.
The results of the proposed Transformer model are presented in Table V. In addition, the DM tests comparing the Transformer against the other benchmark models are also shown. Transformer model has significantly better results (DM test under 0.05) than the model from [2] in four of the five datasets. It also has significantly better results than the naive model in all the datasets. The results evidence the improvement in adding the transformer architecture to the price time series.
@@ IV Results
TABLE II: Naive model test results
@@
TABLE III: DNN Ensemble renormalized with all past values test results
@@
TABLE IV: DNN Ensemble renormalized with last-year values test results
@@
TABLE V: Transformer model test results and DM p-values obtained.
@@
V Conclusion and Future work
Even though the Transformer has been used for forecasting in other domains, there was a lack of conclusive evidence that indicated the performance against other approaches for electricity price forecasting. In this paper, a transformer-based forecasting model is proposed and tested in the framework proposed in [2]. It has been shown that the model offers significant better performance for the majority of the cases, having state-of-the-art results for four of the five datasets. Hence, the Transformer architecture offers good prediction results for electricity price forecasting.
Regarding future work, there could be several lines of research worth exploring. For example, the authors in [2] claim that LSTMs could potentially be more accurate than DNN, but the evidence was insufficient. A possible future study would be a comparison between LSTMs and Transformers. It is important to note that the model presented in this paper is not an Ensemble. An Ensemble of Transformers could have been better, but it was out of the scope of this paper. However, it would be a relevant field of study. Finally, another future work would be to explore how to take also into account past values of the exogenous variables in the transformer arquitecture.

### v3

# inst1]organization=Ericsson Cognitive Labs, addressline=Retama Ed 1 Torre Suecia, city=Madrid, postcode=28045, state=Madrid, country=Spain

@@ Abstract
keywords:
[…]
1 Introduction
@@
Forecasting electricity prices plays a pivotal role in the modern power systems landscape. Accurate and efficient price forecasting methods have become increasingly crucial as the electricity market continues to evolve towards a more competitive and deregulated structure  Weron (2014); Lago et al. (2021) . These forecasts serve as a fundamental decision-making tool for various market participants, including power producers, consumers, traders, and grid operators.
@@
Furthermore, with the growing integration of renewable energy sources, which are inherently intermittent and unpredictable, the volatility and complexity of electricity prices have increased. This motivates the need for more sophisticated and robust forecasting methods to capture these complex dynamics and provide accurate price forecasts  Bhatia et al. (2021). Thus, developing effective electricity price forecasting methods is not only an economic imperative but also a key enabler for the sustainable and reliable operation of power systems.
@@
The techniques for electricity price forecasting (EPF) have evolved significantly over the years, with recent advancements in machine learning and artificial intelligence leading the way  Zhang et al. (2020). While effective in certain scenarios, traditional statistical methods often struggle to capture the complex, non-linear dynamics of electricity prices. This has led to the exploration of more sophisticated models, particularly neural networks and deep learning models applied to time series forecasting. For example, Recurrent Neural Networks (RNNs) and Long Short Term Memory (LSTM) networks have been successfully employed for EPF  Memarzadeh and Keynia (2021). These models are capable of learning complex patterns and long-term dependencies, making them well-suited for forecasting tasks.
@@
Among the deep learning models, the application of Transformer-based architecture is an emerging technique. The design of attention mechanisms for approaching sequence-to-sequence problems started in the NLP field, where the output is not only a label, but a multiple sequence (i.e. a complete sentence). Traditionally an encoder-decoder structure was used until the Attention layer was presented Bahdanau et al. (2016). The Attention layer allows the decoder to focus its attention on a specific word or group of words from the input of the encoder (the original sentence in the case of translation). This helped improve the State of the Art of many NLP problems, becoming the basis of the Transformer model. One of the advantages of the Transformer architecture is that it enables the generation of much bigger Deep Learning models (such as the BERT model Devlin et al. (2019a)).
@@
The attention mechanisms that allow focus on the most relevant parts of the input make them powerful tools for prediction tasks and have been successfully used in several fields such as Autonomous Vehicles Chib and Singh (2024) or Image Classification Dosovitskiy et al. (2022).
@@
Attention mechanisms have also been applied to the power sector. For example,  Marulanda et al. (2023) use attention mechanisms combined with LSTM models for short-term wind power forecasting. Other examples include Zhao et al. (2021) and Chen et al. (2019) which propose the use of attention and transformers for electricity load forecasting.
@@
Regarding electricity price forecasting, additional challenges are presented as input explanatory variables have to be included as relevant predictors in the model, hence, Bottieau et al. (2023) uses an attention layer combined with BLSTM layers for electricity price prediction to achieve promising results. Additionally, Li et al. (2022) propose a Transformer-based model combined with GRU layers for probabilistic electricity price forecasting.
@@
Therefore, the application of Transformer models in forecasting is still in its early stages, and further research is needed to understand their potential and limitations with respect to other forecasting methods. In fact, most of the mentioned models were not Transformers, but a recurrent model combined with attention layers. Authors such as Zeng et al. (2023), still question the effectiveness of Transformer capabilities, hence, making this a relevant issue in the research community.
@@
Additionally, in the EPF field, the issue of reproducibility has been a significant challenge. Many studies have utilized unique, non-public datasets and have tested their methods over too short and limited market samples, making it difficult to evaluate the effectiveness of new predictive algorithms. In response to these issues,  Lago et al. (2021) proposed the EPF toolbox, an open-source toolbox for electricity price forecasting that enhances reproducibility, promotes transparency and drives innovation.
@@
2 Proposed Transformer model for Electricity Price Forecasting
This section explains the proposed model for EPF. The objective is the same as stated in Lago et al. (2021): to predict the electricity prices of the 24h of the day $D+1$ based on the exogenous variables of the 24h of the day $D+1$ as well as past prices and exogenous variables.
The model proposed in this paper is a pure Transformer, specifically, a Transformer Encoder as in the case of BERT Devlin et al. (2019b) or ViT Dosovitskiy et al. (2022). With this structure, the Transformer can capture temporal patterns and make a prediction focusing more attention on longer trends than other temporal models (e.g. LSTMs). This is evidenced in the case of Natural Language Processing  Vaswani et al. (2017), where the Transformer does not forget the previous information in longer sequences.
@@ 2 Proposed Transformer model for Electricity Price Forecasting
We ought to establish a direct relationship between exogenous variables and prices and keep the Transformer to model the price dynamics. Therefore, it is decided to include only past prices as historical information. Therefore, this model will have two inputs, the exogenous values for the 24h of the day $D+1$ and the prices of the 24h for the past $m$ days. The general structure can be visualized in Figure 1. Then, the interval used for the past prices will be [DmD-m, $D$], including both days. These $m$ days will be a hyperparameter of the model called sequence length that should be optimized.
@@
3 Case Study
As pointed out in Lago et al. (2021), in the literature the comparisons between different types of models have been unfair most times. They detected the following problems in the published literature on EPF models:
@@ 3 Case Study
Hence they proposed a new framework for testing the different models: epftoolbox Wel . This python library allows to use the models they have developed and also test the different datasets to build a fair comparison, solving all the issues that have been mentioned previously. Therefore, in this paper, this framework is used to compare and test the model presented.
3.1 Open Access Datasets
To find the best algorithm for electricity price forecasting, it is important to consider multiple datasets, rather than relying on a single one. In Lago et al. (2021), five different datasets for different countries are available so methods can be compared fairly:
@@ 3.1 Open Access Datasets
3.2 Train-Validation-Test split
The testing period is a source of discussion. Most articles try to evaluate algorithms in short periods, e.g. a week or a month. However, the dynamics in a market can change throughout the year, especially in the electricity domain, which is highly affected by the time of the year due to issues such as the available amount of renewable energies or the time consumers are spending at home (great difference between summer and winter). Following Lago et al. (2021), a two-year testing period is used to account the different dynamics that can appear.
@@ 3.2 Train-Validation-Test split
Table 1: Train-Test split
3.3 Retraining and Renormalization
In Lago et al. (2021), the authors state that a model should be retrained instead of being directly evaluated in the whole dataset, or otherwise it is not tested in realistic conditions, since in real-life forecasting models are often retrained to account for the last market information. Even though this idea is logical, it generates two main problems for testing more advanced models such as the one proposed in this paper
On the one hand, although retraining each day simulate well real life behaviour, there are some difficulties to automatize that retraining for more complicated models. While retraining a Deep Neural Network composed of linear layers and activation functions seems to follow a straightforward process explained in Lago et al. (2021), where early-stopping gives the optimal training point for the selected hyperparameters, more complicated models as Transformers can be more difficult to train due to issues like vanishing or exploding gradients or the amplification effect pointed put in Liu et al. (2020). Therefore for these models, it is more likely that a Machine Learning Engineer monitors the retraining phase and execute it if necessary. It could be the case that for some time the parameters for the retraining are the same, and only when the data changes are significant enough it is necessary to look again for optimal values.
On the other hand, the computational burden is expensive for advanced models. While the best model in Lago et al. (2021) takes around five minutes to train, the Transformer model presented in this paper takes around twenty minutes (measured in a computer with a GPU RTX 3060).
@@ 3.3 Retraining and Renormalization
Nevertheless, as suggested in Lago et al. (2021), a renormalization step was applied consisting on subtracting the mean and standard deviation from the data previous to the next day to predict. This helps the model to adapt if the conditions vary greatly during the test period.
3.4 Models to Compare
@@
  • DNN: Regarding neural networks in  Lago et al. (2021), a Deep Neural Network with two layers was selected as the best Machine Learning Model. In fact, the model that proved to be more accurate in their study was the Ensemble of Deep Neural Networks. Therefore, this model ensemble is selected as a benchmark. As it was explained the ensemble will be trained on the training period and evaluated during the test period using a renormalization step but not retraining. We compute two approaches, a renormalization with all the past data and a renormalization with only the last year values (as in Lago et al. (2021)).
@@
  • Transformer model: This is the transformer model proposed. As mentioned in Section 2, some hyperparameters have to be optimized using a validation set. Following what was proposed in Lago et al. (2021), validation is constructed with the last 42 weeks of the training set. Some values were proposed for the different hyperparameters and datasets and the configuration with the best performance is selected. In Tables 6, 7, 8, 9 and 10 the results of some of the configurations in the different datasets are presented along with the best ones highlighted in gray. The best model is evaluated then on the test set, without any retraining but with data renormalization with all past values.
Following what was exposed in Lago et al. (2021), to measure the results offered by the model presented in this paper and compare it with the benchmarks explained, the following metrics will be used: MAE, RMSE and sMAPE.
3.5 Statistical Testing
@@
In particular, the one-sided DM test is implemented in Lago et al. (2021) and used in the case study. The null hypothesis of the test is that the predictions of model A are more accurate than the ones of a model B. The lower the p-value obtained, the more observed data inconsistent with the null hypothesis. Taking into account the convention to reject the null hypothesis at p-values lower than 5%, if a lower p-value is obtained, it can be declared that the difference between predictions of model B and predictions of model A are statistically significant, being B the model with better performance.
3.6 Software
The aim of this paper is to compare the Transformer model presented with the ones developed in the  Lago et al. (2021). Hence, every implementation from the mentioned library will be used in order to follow a rigorous approach to compare the different methods. As the aim of this paper is to openly contribute to EPF research, all the code and the best models can be found in https://github.com/osllogon/epf_transformers.
4 Results
The results of the selected models tested on the different datasets are presented in this section. For the comparison, the different metrics explained in the Section 3.4 are calculated for the test set predictions.
Tables 2, 3 and 4 show the benchmark Naïve, DNN with all past renormalization and DNN with last-year renormalization model results, respectively. As can be observed, in some cases the results of the DNN model are better and, in other cases, the naive model outperforms the DNN Ensembles. This could be due to the fact that the DNN model has to make predictions for a long test period without retraining.
The results of the proposed Transformer model are presented in Table 5. In addition, the DM tests comparing the Transformer against the other benchmark models are also shown. Transformer model has significantly better results (DM test under 0.05) than the model from Lago et al. (2021) in four of the five datasets. It also has significantly better results than the naive model in all the datasets. The results evidence the improvement in adding the transformer architecture to the price time series.
Table 2: Naive model test results
@@ 4 Results
Table 3: DNN Ensemble renormalized with all past values test results
@@
Table 4: DNN Ensemble renormalized with last-year values test results
@@
Table 5: Transformer model test results and DM p-values obtained.
@@
5 Conclusion and Future work
Even though the Transformer has been used for forecasting in other domains, there was a lack of conclusive evidence that indicated the performance against other approaches for electricity price forecasting. In this paper, a transformer-based forecasting model is proposed and tested in the framework proposed in Lago et al. (2021). It has been shown that the model offers significant better performance for the majority of the cases, having state-of-the-art results for four of the five datasets. Hence, the Transformer architecture offers good prediction results for electricity price forecasting.
Regarding future work, there could be several lines of research worth exploring. For example, the authors in Lago et al. (2021) claim that LSTMs could potentially be more accurate than DNN, but the evidence was insufficient. A possible future study would be a comparison between LSTMs and Transformers. It is important to note that the model presented in this paper is not an Ensemble. An Ensemble of Transformers could have been better, but it was out of the scope of this paper. However, it would be a relevant field of study. Finally, another future work would be to explore how to take also into account past values of the exogenous variables in the transformer arquitecture.
Appendix A Validation results for the proposed model arquitecture
Table 6: Nord Pool Validation Results
$n$ NH NL FD P Opt SL Lr Sch CG MAE
128 1 6 1024 0.2 Adam 672 1e-4 70-0.1 None 1.336
512 8 6 2048 0.2 Adam 672 1e-5 70-0.1 None 1.264
512 8 6 1024 0.2 Adam 336 1e-5 70-0.1 None 1.196
256 8 6 1024 0.1 Adam 672 1e-5 70-0.1 None 1.225
128 8 6 512 0.1 Adam 672 1e-5 70-0.1 None 1.243
Table 7: PJM Validation Results
$n$ NH NL FD P Opt SL Lr Sch CG MAE
128 4 4 1048 0.2 Adam 336 1e-4 70-0.1 0.1 2.526
64 2 4 256 0.2 Adam 336 1e-4 70-0.1 0.1 2.576
64 4 4 1024 0.2 Adam 336 1e-4 None 0.1 2.626
64 4 4 1024 0.2 Adam 336 1e-3 70-0.1 0.1 2.699
64 8 6 1024 0.2 Adam 336 1e-5 None 0.1 2.811
Table 8: EPEX-FR Validation Results
$n$ NH NL FD P Opt SL Lr Sch CG MAE
512 1 6 1024 0.2 Adam 672 1e-5 50-0.2 0.25 3.778
512 1 6 2048 0.2 Adam 672 1e-5 50-0.2 0.25 3.853
512 8 6 1024 0.2 Adam 336 1e-5 50-0.2 0.25 3.756
512 8 6 1024 0.2 Adam 672 1e-5 50-0.2 0.25 3.732
512 8 6 1024 0.2 Adam 672 1e-5 70-0.2 0.25 3.751
512 8 6 102 0.2 Adam 672 1e-5 70-0.1 0.5 3.759
Table 9: EPEX-DE Validation Results
$n$ NH NL FD P Opt SL Lr Sch CG MAE
256 4 4 1024 0.2 Adam 336 1e-4 None 0.5 3.904
256 8 6 1024 0.2 Adam 336 1e-5 None 0.25 3.577
512 8 6 2048 0.2 Adam 672 1e-5 None 0.25 3.886
512 8 6 2048 0.2 Adam 672 1e-6 None 0.25 3.821
64 4 4 256 0.2 Adam 336 1e-5 None 0.5 3.887
64 8 6 256 0.2 Adam 336 1e-5 None 0.25 3.873
Table 10: EPEX-BE Validation Results
$n$ NH NL FD P Opt SL Lr Sch CG MAE
128 4 4 1024 0.2 Adam 672 1e-5 None 0.25 5.238
128 8 6 1024 0.2 Adam 672 1e-5 None 0.25 5.265
64 4 4 1024 0.2 Adam 672 1e-5 None 0.5 5.324
64 4 4 256 0.2 Adam 336 1e-4 None 0.1 5.125
64 4 4 256 0.2 Adam 336 1e-5 None 0.1 4.903
64 4 4 256 0.2 Adam 672 1e-5 None 0.1 4.854